# DRL-Based Federated Self-Supervised Learning for Task Offloading and Resource Allocation in ISAC-Enabled Vehicle Edge Computing

### v1

@@ Abstract
Integrated Sensing and Communications (ISAC), Federated Self-Supervised Learning, Resource Allocation and Offloading, Deep Reinforcement Learning (DRL), Vehicle Edge Computing
footnotetext: Email addresses: xueyinggu@stu.jiangnan.edu.cn(X. Gu), fpy@tsinghua.edu.cn(P. Fan), dr.nan.cheng@ieee.org(N. Cheng), wenchen@sjtu.edu.cn(W. Chen), eekhaled@ust.hk(K. B. Letaief)footnotetext: This work was supported in part by the National Natural Science Foundation of China under Grant No. 61701197, in part by the National Key Research and Development Program of China under Grant No.2021YFA1000500(4), in part by the Research Grants Council under the Areas of Excellence scheme grant AoE/E-601/22-R, and in part by the 111 Project under Grant No. B12018.
@@ Introduction
Intelligent Transportation Systems (ITS) capitalize on the integration of sensing and communication technologies, known as Integrated Sensing and Communications (ISAC), to significantly improve the efficiency of data exchange between vehicles and infrastructure within the Internet of Vehicles (IoV) [1, 2, 3, 4]. However, this also brings about significant computing demands that often exceed local computing capability, potentially hindering real-time data processing [5, 6, 7, 8]. Vehicle Edge Computing (VEC) addresses this issue by deploying computing resource at the network edge, specifically at Road Side Unit (RSU) [9, 10, 11, 12]. As integral components of the vehicular communication infrastructure, RSU is equipped with computing power and storage capacity, thereby delivering real-time computing services and ensuring timely task completion [13, 14, 15, 16]. Meanwhile, offloading partial computing task to RSU can significantly reduce the computing burden on moving vehicles in IoV.
@@
We previously proposed a Self-Supervised Learning (SSL) method based on Federated Learning (FL) called FLSimCo algorithm [17]. In each round of FL, each vehicle can calculate the expected number of local training iterations for SSL based on utilizing the entire computing resource of vehicle. However, for efficiency and practicality, vehicle often has multiple local tasks that need to process within a fixed time, including system maintenance, running other applications, and background processes [18, 19, 20, 21]. Due to CPU usage by other tasks, the expected local training iterations are often hard to fully execute in practice. In this case, offloading partial training iterations task to RSU can ensure the expected local training iterations task is completed within one round of FL [22, 23, 24, 25].
@@
Compared to the vast computing resource of centralized cloud server, if the resource of RSU is not allocated properly, it will be difficult to meet the numerous task offloading requests from vehicles during peak periods[26, 27, 28, 29]. A simple solution is to deploy more RSUs, but this significantly increases costs of info-structure and may lead to substantial resource waste during off-peak periods. Therefore, this paper aims to maximize the utilization of a single RSU in a four-way roundabout, rationally allocate its computing resource, and meet more vehicle offloading requests [30, 31, 32, 33].
@@
In this paper, we further explore the potential of efficient offloading decisions based on the FLSimCo algorithm in IoV. We propose an improved algorithm that BS employs SAC algorithm to allocate transmission powers from vehicles to BS and RSU, CPU frequencies in vehicles, assignment ratios for computing resource in RSU. According to the allocated assignment ratio, the computing resource in RSU allowed for each vehicle can be calculated. Then according to the allowed computing resource, the local training iterations task in each round of FLSimCo is divided into local and RSU assisted training. However, at lower assignment ratio, it suggests the allocated computing resource in RSU is small. In this case, if vehicle offloads task to RSU, it does not ease the computing burden on vehicle and also increases communication costs on the requested tasks. Therefore, we also set an offloading threshold to optimize offloading efficiency. After offloading partial iterations task to RSU, the vehicle processes the remaining iterations task locally. If there are still unfinished iterations in round of FL, the remaining number of iterations will be stored in local buffer of vehicle, awaiting for the next training [^1] The source code has been released at: https://github.com/qiongwu86/Federated-SSL-task-offloading-and-resource-allocation.
@@
In IoV, task offloading and computing resource allocation have been longstanding research hot spots [34, 35, 36, 37, 38]. With the increasing number of vehicles and the rising quantity and complexity of computing tasks, efficiently allocating and managing resource have become a focal point of attention. In this section, we will introduce the related work about task offloading and resource allocation.
@@
With the rapid development of machine learning, an increasing number of work focus on applying machine learning to task offloading and resource allocation [39, 40, 41, 42]. Due to its superior performance in handling complex dynamic environments, DRL has gradually become a popular choice for optimizing task offloading and resource allocation.
@@
Some work combined FL and SSL [51][52], but these methods mostly based on static environments, relying heavily on extensive negative samples during the SSL process, which leads to extended training duration. However, in real-time demanding environment such as IoV, constraints related to time and environmental variability become significant. Therefore, conducting SSL in IoV necessitates addressing the challenge of reducing dependency on negative samples during the training process.
@@
In this section, we will introduce the system scenario and models in this paper. As shown in Fig. 1, the scenario is a VEC traffic system containing four intersections. In this scenario, a BS is located at the corner, an RSU is deployed by the roadside, and $N$ vehicles are moving on the roads. We assume that all four intersections are within the coverage area of both BS and RSU. When vehicle $n$, $n\in[1,N]$, arrives at an intersection, it chooses to turn left, turn right, or go straight through the intersection with probabilities of $\psi_{1}$, $\psi_{2}$, and $\psi_{3}$, respectively. The neural network configured on the BS is used for decision support. The RSU has fixed computing resource to parallel process tasks offloaded from vehicles.
@@
The entire process consists of two stages: offloading-based Federated SSL and DRL-based resource allocation. These two stages are interdependent yet mutually influential. The stage of offload-based Federated SSL consists $E^{max}$ episodes, and each episode $e$, $e\in[1,E^{max}]$, is divided into $R^{max}$ rounds. Similarly, the stage of DRL-based resource allocation consists $E^{max}$ episodes, and each episode $e$ is divided into $S^{max}$ slots. Each round of FL corresponds to a slot, and the duration is $T$. However, when slot $t$ reaches to $S^{max}$, the episode $e$ enter to the next episode $e+1$, and slot $t$ resets to zero. Meanwhile, the round $r$ continues to accumulate until the entire algorithm is completed. Thus, the relationship between round $r$ and slot $t$ in episode $e$ is described as
$r=\left(e-1\right)\cdot S^{max}+t,\quad e\in[1,E^{max}]\text{ and }t\in[1,S^{max}].$ (1)
Before the start of each slot $t$, the BS stores the global model aggregated from round $r-1$ for FL and uses the DRL to allocate resource, including the transmission power $p_{t,R}^{n}$ and $p_{t,B}^{n}$ for transmitting data to the RSU and BS, the CPU frequency $f_{t,L}^{n,comp}$ for local training, and assignment ratio $q_{t}^{n}$ for vehicle $n$. When episode $e=1$, if in the first round/slot, BS randomly initializes the parameters of global model and DRL network. Meanwhile, vehicle $n$ stores $\mathcal{Z}$ images captured during previous round, denoted as local data, and a model, denoted as the local model, that has the same structure as the global model in BS side. In each slot $t$, vehicle $n$ first downloads the data stored in the BS and assign the parameters of global model from BS to their local model. Next, by the downloaded CPU frequency $f_{t,L}^{n,comp}$, the theoretical total number of local training iterations $N_{t}^{n}$ for vehicle $n$ is calculated, and the actual total number of training iterations ${N_{t}^{n}}^{\ast}$ for vehicle $n$ also can be calculated based on the available CPU time $T^{\prime}$ of vehicle $n$. Finally, vehicle $n$ decides whether to offload the computation tasks to the RSU in slot $t$ based on the downloaded assignment ratio $q_{t}^{n}$ and the assignment ratio threshold $q_{0}$. Specifically, when $q_{t}^{n}$ is greater than $q_{0}$, the vehicle $n$ will offload part of the tasks to RSU; otherwise, no task will be offloaded, and $q_{t}^{n}$ will be reset to 00. If decides to offload task, the vehicle $n$ immediately sends an offloading request to RSU with the downloaded transmission power $p_{t,R}^{n}$. The request includes the local model, local data, and the assignment ratio of vehicle $n$ in slot $t$.
After receiving the offloading request from vehicles, the RSU allocates computing resource based on its own CPU frequency $f_{t,R}^{comp}$ and the received assignment ratios $\left\{q_{t}^{1},q_{t}^{2},\cdots q_{t}^{n},\cdots q_{t}^{N}\right\}$. Meanwhile, vehicle $n$ also utilizes local model and local data to complete the local training of the non-offloaded portion. Once the RSU completes all training tasks, it sends back the trained model to each vehicle. Upon receiving the model trained by the RSU, vehicle $n$ uploads both models trained by locally and RSU to BS with the downloaded transmission power $p_{t,B}^{n}$, where the models will be aggregated. Next, we will introduce the computation model, channel model, and transmission model for vehicle $n$ in slot $t$.
@@
During the training process, the energy consumed for performing one training iteration on vehicle $n$ is calculated as follows:
$E_{t,L}^{n,comp}=p_{t,L}^{n,comp}T_{t,L}^{n,comp},$ (2)
where $p_{t,L}^{n,comp}$, $T_{t,L}^{n,comp}$, and $E_{t,L}^{n,comp}$ represent the computing power, delay, and energy consumption, respectively, for vehicle $n$ to perform one local training iteration. The local CPU computation power $p_{t,L}^{n,comp}$ of vehicle $n$ in slot $t$ can be calculated using the Dynamic Voltage and Frequency Scaling (DVFS) method as follows:
$p_{t,L}^{n,comp}=\kappa\left(f_{t,L}^{n,comp}\right)^{3},$ (3)
@@
To calculate the computation delay of vehicle $n$, let $c$ be the number of CPU cycles required to process a unit size of data. $Z$ is the size of the local data, and it is assumed that each vehicle has the same data size. Then, $cZ$ represents the number of CPU cycles needed for one iteration. Therefore, the computation delay $T_{t,L}^{n,comp}$ is calculated as follows:
$T_{t,L}^{n,comp}=\frac{cZ}{f_{t,L}^{n,comp}}.$ (4)
By substituting Eqs. (3) and (4) into Eq. (2), we obtain:
$E_{t,L}^{n,comp\ }=\kappa\left.\left(f_{t,L}^{n,comp\ }\right.\right)^{2}cZ.$ (5)
@@
$T_{t,R}^{comp}=\frac{cZ}{f_{t,R}^{comp}},$ (6)
$E_{t,R}^{comp\ }=\kappa\left.\left(f_{t,R}^{comp\ }\right.\right)^{2}cZ,$ (7)
@@
The vehicles involved in training use Orthogonal Frequency Division Multiple (OFDM) technology to communicate with the RSU [58], so vehicle-to-vehicle interference is not considered during communication. Each vehicle uses the same channel model in the same slot, and the channel model is updated as the slot $t$ advances to slot $t+1$. Therefore, in slot $t$, the channel gain can be calculated as:
$h_{t,R}^{n,trans}\left[\text{dis}(R,n)\right]=\alpha_{t}^{n}\left[\text{dis}(R,n)\right]m_{t}^{n},$ (8)
where $m_{t}^{n}$ denotes small-scale fading, which follows an exponential distribution with a unit mean, i.e., $m_{t}^{n}\sim\text{Exp}(1)$. $\text{dis}(R,n)$ represents the distance between vehicle $n$ and the RSU, and measured in meters. $\alpha_{t}^{n}$ represents the effect of large-scale fading on the signal, including path loss $\mathcal{P}$ and shadow fading $\mathcal{S}$. Thus, $\alpha_{t}^{n}$ can be calculated as:
$\alpha_{t}^{n}={10}^{\frac{\mathcal{P}}{10}}{10}^{\frac{\mathcal{S}}{10}},$ (9)
where shadow fading $\mathcal{S}$ follows a normal distribution with a mean of $\mu$ and a standard deviation of $\sigma$, represented as $\mathcal{S}\sim\mathcal{N}\left(\mu,\sigma^{2}\right)$. In high-density urban environments, the path loss $\mathcal{P}$ can be calculated as:
$\mathcal{P}=128.1+37.6\log\left[\frac{\text{dis}(R,n)}{1000}\right].$ (10)
Therefore, the information transmission rate $R_{t,R}^{n}$ between the vehicle $n$ and the RSU can be calculated as [59]:
$R_{t,R}^{n}=B^{n}\log_{2}\left[1+\frac{p_{t,R}^{n}h_{t,R}^{n,trans}\text{dis}(R,n)}{N_{0}}\right],$ (11)
@@
If vehicle $n$ offloads partial iterations task to the RSU, during the transmission process, the energy consumption $E_{t}^{n,trans}$ can be calculated as:
$E_{t}^{n,trans}=p_{t,R}^{n}T_{t,R}^{n,trans},$ (12)
where $T_{t,R}^{n,trans}$ is the time required for vehicle $n$ to transmit data to RSU. Assuming each vehicle transmits a local model of the same size $D$, and the data transmitted by vehicle $n$ to RSU includes the untrained local model and local data. Therefore, the transmission delay $T_{t,R}^{n,trans}$ can be calculated as:
$T_{t,R}^{n,trans}=\frac{Z+D}{R_{t,R}^{n}}.$ (13)
@@
When vehicles are performing local iterations task, if they are unable to complete all iterations locally due to local computing resource limitations, partial iterations task needs to be offloaded to RSU. This offloading process requires BS to use a DRL algorithm to allocate assignment ratios, which determines the computing resource of the RSU allowed for each vehicle. The total iterations are then allocated between local training and training on the RSU. This optimization aims to enhance resource utilization and task completion efficiency. Next, we will delve into the specific methods of task allocation.
@@
$N_{t}^{n}=\frac{T-t^{max}}{T_{t,L}^{n,comp}},$ (15)
where $t^{max}$ represents the maximum time used for data transmission. Thus, $T-t^{max}$ indicates the total time available for computing within a slot. In practical, the CPU typically handles multiple tasks simultaneously to improve CPU utilization. These tasks consume some CPU computing resources, resulting in the actual time available for local iterations being less than $T$. We denote the true training time as $T^{\prime}$. Therefore, the actual total local training iterations ${N_{t}^{n}}^{\ast}$ can be calculated as:
${N_{t}^{n}}^{\ast}=\frac{T^{\prime}-t^{max}}{T_{t,L}^{n,comp}}.$ (16)
Substitute Eq. (4) into Eq. (15) and Eq. (16), the final results can be obtained, respectively:
$N_{t}^{n}=\frac{(T-t^{max})f_{t,L}^{n,comp}}{cZ},$ (17)
${N_{t}^{n}}^{\ast}=\frac{(T^{\prime}-t^{max})f_{t,L}^{n,comp}}{cZ}.$ (18)
In this paper, we study a binary partial offloading problem, where a binary variable $g_{t}^{n}$ is used to indicate whether vehicle $n$ needs to offload partial task to the RSU. $g_{t}^{n}$ is related to the assignment ratio $q_{t}^{n}$ and the assignment ratio threshold $q_{0}$. When $q_{t}^{n}$ is less than the assignment ratio threshold $q_{0}$, it indicates that vehicle $n$ will not offload partial task to RSU and reset $q_{t}^{n}=0$; otherwise, it will offload it. Therefore, $g_{t}^{n}$ can be expressed as:
$g_{t}^{n}=\begin{cases}1&\text{if }q_{t}^{n}\geq q_{0}\\ 0&\text{if }q_{t}^{n}<q_{0}\end{cases}.$ (19)
@@
We denote the total task that vehicle $n$ needs to handle in slot $t$ as $N_{t}^{n,total}$, which includes the sum of the total training iterations $N_{t}^{n}$ in slot $t$ and the remaining unprocessed training iterations $B_{t-1}^{n}$ from the previous slot. When vehicle $n$ offloads partial task to RSU, the total task $N_{t}^{n,total}$ will be divided into local training iterations $N_{t,L}^{n}$ and offloaded training iterations $N_{t,R}^{n,off}$ according to the allocated assignment ratio $q_{t}^{n}$. Specifically, RSU first determines the total training iterations $N_{R}$ in RSU side as follows:
$N_{R}=\frac{T-t^{max}-T_{t,R}^{n,trans}}{T_{t,R}^{comp}}.$ (20)
Thus, the expected offloaded training iterations $N_{t,R}^{n,off}$ for RSU can be calculated as:
$N_{t,R}^{n,off}=g_{t}^{n}q_{t}^{n}N_{R}.$ (21)
@@
$N_{t,R}^{n,off}\leftarrow\min\left(N_{t,R}^{n,off},N_{t}^{n,total}\right).$ (22)
@@
$B_{t}^{n}=\begin{cases}0&\text{if}N_{t}^{n,re}\leq{N_{t}^{n}}^{\ast}\\ N_{t}^{n,re}-{N_{t}^{n}}^{\ast}&\text{if }N_{t}^{n,re}>{N_{t}^{n}}^{\ast}\end{cases}.$ (25)
@@
To achieve efficient task offloading and resource allocation, it is essential for BS to consider the long-term impact of their actions during the decision-making process. Therefore, we aim to minimize the sum of the energy consumption of all vehicles in each slot. Let =$\mathbb{P}=\left\{p_{t,R}^{1},p_{t,R}^{2},\cdots p_{t,R}^{n},\cdots p_{t,R}^{N}\right\}$, $\mathbb{F}=\left\{f_{t,L}^{1,comp},f_{t,L}^{2,comp},\cdots f_{t,L}^{n,comp},\cdots f_{t,L}^{N,comp}\right\}$ and =$\mathbb{Q}=\left\{q_{t}^{1},q_{t}^{2},\cdots q_{t}^{n},\cdots q_{t}^{N}\right\}$. Then, we can formulate the optimization problem in slot $t$ as follows:
@@
$g_{t}^{n}\in\left\{0,1\right\},\forall n\in N$ (29a)
$p^{min}\leq p_{t,R}^{n}\leq p^{max},\forall n\in N$ (29b)
$f^{min}\leq f_{t,L}^{n,comp}\leq f^{max},\forall n\in N$ (29c)
$0\leq q_{t}^{n}\leq 1,\forall n\in N$ (29d)
@@
After the BS allocates the assignment ratio $q_{t}^{n}$ for vehicle $n$ in slot $t$, the iterations number of offloading task $N_{t}^{n,off}$ assigned to vehicle $n$ might exceed the total number of tasks $N_{t}^{n,total}$ that the vehicle $n$ needs to process in slot $t$ [44]. We denote the excess part as the overload $O_{t}^{n}$, which can be calculated as:
$O_{t}^{n}=\max\left(0,N_{t,R}^{n,off}-N_{t}^{n,total}\right).$ (31)
@@
$\mathcal{R}_{0}=\frac{1}{NS^{max}}\left(\sum_{n=1}^{N}\sum_{t=1}^{S^{max}}{\frac{O_{t}^{n}}{N_{t}^{n,total}}}\right).$ (32)

### v2

@@ Abstract
Integrated sensing and communications (ISAC), Federated self-supervised learning, Resource allocation and offloading, Deep reinforcement learning (DRL), Vehicle edge computing (VEC)
footnotetext: Email addresses: xueyinggu@stu.jiangnan.edu.cn(X. Gu), fpy@tsinghua.edu.cn(P. Fan), dr.nan.cheng@ieee.org(N. Cheng), wenchen@sjtu.edu.cn(W. Chen), eekhaled@ust.hk(K. B. Letaief)
@@ Introduction
As Intelligent Transportation Systems (ITS) evolve, they increasingly rely on the integration of sensing and communication technologies, collectively known as Integrated Sensing and Communications (ISAC), to significantly improve the efficiency of data exchange between vehicles and infrastructure within the Internet of Vehicles (IoV) [1, 2]. However, this also brings about significant computing demands that often exceed local computing capability, potentially hindering real-time data processing [3, 4, 5]. To address these escalating computing demands, Vehicle Edge Computing (VEC) has emerged as a viable solution by deploying computing resources at the network edge, specifically at Road Side Units (RSUs) [6, 7, 8]. As integral components of the vehicular communication infrastructure, RSU is equipped with computing power and storage capacity, thereby delivering real-time computing services and ensuring timely task completion [9, 10, 11]. Meanwhile, offloading partial computing task to RSU can significantly reduce the computing burden on moving vehicles in IoV.
@@
We previously proposed a Self-Supervised Learning (SSL) method based on Federated Learning (FL) called FLSimCo algorithm [12]. In each round of FL, each vehicle can calculate the expected number of local training iterations for SSL based on utilizing the entire computing resource of vehicle. However, for efficiency and practicality, vehicle often has multiple local tasks that need to process within a fixed time, including system maintenance, running other applications, and background processes [13, 14, 15]. This need for local task processing introduces competition for computing resources, complicating the execution of expected local training iterations. Due to CPU usage by other tasks, the expected local training iterations are often hard to fully execute in practice. In this case, offloading partial training iterations task to RSU can ensure the expected local training iterations task is completed within one round of FL [16, 17].
@@
Compared to the extensive computing resources available in centralized cloud servers, the effective allocation of RSU resources becomes critical to meeting the demands of numerous task offloading requests from vehicles, particularly during peak periods [18, 19]. A simple solution is to deploy more RSUs, but this significantly increases costs of info-structure and may lead to substantial resource waste during off-peak periods. Therefore, this paper aims to maximize the utilization of a single RSU in a four-way roundabout, rationally allocate its computing resource, and meet more vehicle offloading requests [20, 21].
@@
In this paper, we further explore the potential of efficient offloading decisions based on the FLSimCo algorithm in IoV. We propose an improved algorithm that BS employs SAC algorithm to allocate transmission powers from vehicles to BS and RSU, CPU frequencies in vehicles, assignment ratios for computing resource in RSU. According to the allocated assignment ratio, the computing resource in RSU allowed for each vehicle can be calculated. Then according to the allowed computing resource, the local training iterations task in each round of FLSimCo is divided into local and RSU assisted training. However, at lower assignment ratio, it suggests the allocated computing resource in RSU is small. In this case, if vehicle offloads task to RSU, it does not ease the computing burden on vehicle and also increases communication costs on the requested tasks. Therefore, we also set an offloading threshold to optimize offloading efficiency. After offloading partial iterations task to RSU, the vehicle processes the remaining iterations task locally. If there are still unfinished iterations in round of FL, the remaining number of iterations will be stored in local buffer of vehicle, awaiting for the next training [^1] The source code has been released at: https://github.com/qiongwu86/Federated-SSL-task-offloading-and-resource-allocation [^2] This work was supported in part by the National Natural Science Foundation of China under Grant 61701197 and Grant 62071296; in part by the National Key Research and Development Program of China under Grant 2021YFA1000500(4); in part by the National Key Project under Grant 2020YFB1807700; in part by Shanghai Kewei under Grant 22JC1404000; in part by the Research Grants Council under the Areas of Excellence Scheme under Grant AoE/E-601/22-R; and in part by the 111 Project under Grant B23008..
@@
In IoV, task offloading and computing resource allocation have been longstanding research hot spots [22, 23, 24]. With the increasing number of vehicles and the rising quantity and complexity of computing tasks, efficiently allocating and managing resource have become a focal point of attention. In this section, we will introduce the related work about task offloading and resource allocation.
@@
With the rapid development of machine learning, an increasing number of work focus on applying machine learning to task offloading and resource allocation [25, 26]. Due to its superior performance in handling complex dynamic environments, DRL has gradually become a popular choice for optimizing task offloading and resource allocation.
@@
Some work combined FL and SSL [35, 36], but these methods mostly based on static environments, relying heavily on extensive negative samples during the SSL process, which leads to extended training duration. However, in real-time demanding environment such as IoV, constraints related to time and environmental variability become significant. Therefore, conducting SSL in IoV necessitates addressing the challenge of reducing dependency on negative samples during the training process.
@@
In this section, we will introduce the system scenario and models in this paper. As shown in Fig. 1, the scenario is an urban VEC traffic system containing four intersections. In this scenario, a BS is located at the corner, an RSU is deployed by the roadside, and $N$ vehicles are moving on the roads. We assume that all four intersections are within the coverage area of both BS and RSU. When vehicle $n$, $n\in[1,N]$, arrives at an intersection, it chooses to turn left, turn right, or go straight through the intersection with probabilities of $\psi_{1}$, $\psi_{2}$, and $\psi_{3}$, respectively. The neural network configured on the BS is used for decision support. The RSU has fixed computing resource to parallel process tasks offloaded from vehicles [41, 42].
@@
The entire process consists of two stages: offloading-based federated SSL and DRL-based resource allocation. These two stages are interdependent yet mutually influential. The stage of offload-based federated SSL consists $E^{max}$ episodes, and each episode $e$, $e\in[1,E^{max}]$, is divided into $R^{max}$ rounds. Similarly, the stage of DRL-based resource allocation also consists $E^{max}$ episodes, and each episode $e$ is divided into $S^{max}$ slots [43, 44, 45]. Each round of FL corresponds to a slot, and the duration is $T$. However, when slot $t$ reaches to $S^{max}$, the episode $e$ enter to the next episode $e+1$, and slot $t$ resets to zero. Meanwhile, the round $r$ continues to accumulate until the entire algorithm is completed. Thus, the relationship between round $r$ and slot $t$ in episode $e$ is described as
$r=\left(e-1\right)\cdot S^{max}+t,\quad e\in[1,E^{max}]\text{ and }t\in[1,S^{max}]$ (1)
Following, we will introduce the computation model, channel model and transmission model in each round/slot.
@@
During the training process, the energy consumed for performing one training iteration on vehicle $n$ is calculated as follows
$E_{t,L}^{n,comp}=p_{t,L}^{n,comp}T_{t,L}^{n,comp}$ (2)
where $p_{t,L}^{n,comp}$, $T_{t,L}^{n,comp}$, and $E_{t,L}^{n,comp}$ represent the computing power, delay, and energy consumption, respectively, for vehicle $n$ to perform one local training iteration. The local CPU computation power $p_{t,L}^{n,comp}$ of vehicle $n$ in slot $t$ can be calculated using the Dynamic Voltage and Frequency Scaling (DVFS) method as follows
$p_{t,L}^{n,comp}=\kappa\left(f_{t,L}^{n,comp}\right)^{3}$ (3)
@@
To calculate the computation delay of vehicle $n$, let $c$ be the number of CPU cycles required to process a unit size of data. $Z$ is the size of the local data, and it is assumed that each vehicle has the same data size. Then, $cZ$ represents the number of CPU cycles needed for one iteration. Therefore, the computation delay $T_{t,L}^{n,comp}$ is calculated as follows
$T_{t,L}^{n,comp}=\frac{cZ}{f_{t,L}^{n,comp}}$ (4)
By substituting Eqs. (3) and (4) into Eq. (2), we obtain
$E_{t,L}^{n,comp\ }=\kappa\left.\left(f_{t,L}^{n,comp\ }\right.\right)^{2}cZ$ (5)
@@
$T_{t,R}^{comp}=\frac{cZ}{f_{t,R}^{comp}}$ (6)
$E_{t,R}^{comp\ }=\kappa\left.\left(f_{t,R}^{comp\ }\right.\right)^{2}cZ$ (7)
@@
The vehicles involved in training use Orthogonal Frequency Division Multiple (OFDM) technology to communicate with the RSU [47, 48], so vehicle-to-vehicle interference is not considered during communication. Each vehicle uses the same channel model in the same slot, and the channel model is updated as the slot $t$ advances to slot $t+1$ [49, 50]. Therefore, in slot $t$, the channel gain can be calculated as
$h_{t,R}^{n,trans}\left[\text{dis}(R,n)\right]=\alpha_{t}^{n}\left[\text{dis}(R,n)\right]m_{t}^{n}$ (8)
where $m_{t}^{n}$ denotes small-scale fading, which follows an exponential distribution with a unit mean, i.e., $m_{t}^{n}\sim\text{Exp}(1)$. $\text{dis}(R,n)$ represents the distance between vehicle $n$ and the RSU, and measured in meters. $\alpha_{t}^{n}$ represents the effect of large-scale fading on the signal, including path loss $\mathcal{P}$ and shadow fading $\mathcal{S}$. Thus, $\alpha_{t}^{n}$ can be calculated as
$\alpha_{t}^{n}={10}^{\frac{\mathcal{P}}{10}}{10}^{\frac{\mathcal{S}}{10}}$ (9)
where shadow fading $\mathcal{S}$ follows a normal distribution with a mean of $\mu$ and a standard deviation of $\sigma$, represented as $\mathcal{S}\sim\mathcal{N}\left(\mu,\sigma^{2}\right)$. In high-density urban environments, the path loss $\mathcal{P}$ can be calculated as
$\mathcal{P}=128.1+37.6\log\left[\frac{\text{dis}(R,n)}{1000}\right]$ (10)
Therefore, the information transmission rate $R_{t,R}^{n}$ between the vehicle $n$ and the RSU can be calculated as [51]
$R_{t,R}^{n}=B^{n}\log_{2}\left[1+\frac{p_{t,R}^{n}h_{t,R}^{n,trans}\text{dis}(R,n)}{N_{0}}\right]$ (11)
@@
If vehicle $n$ offloads partial iterations task to the RSU, during the transmission process, the energy consumption $E_{t}^{n,trans}$ can be calculated as
$E_{t}^{n,trans}=p_{t,R}^{n}T_{t,R}^{n,trans}$ (12)
where $T_{t,R}^{n,trans}$ is the time required for vehicle $n$ to transmit data to RSU. Assuming each vehicle transmits a local model of the same size $D$, and the data transmitted by vehicle $n$ to RSU includes the untrained local model and local data. Therefore, the transmission delay $T_{t,R}^{n,trans}$ can be calculated as
$T_{t,R}^{n,trans}=\frac{Z+D}{R_{t,R}^{n}}$ (13)
@@
When vehicles are performing local iterations task, if they are unable to complete all iterations locally due to local computing resource limitations, partial iterations task needs to be offloaded to RSU [52]. This offloading process requires BS to use a DRL algorithm to allocate assignment ratios, which determines the computing resource of the RSU allowed for each vehicle. The total iterations are then allocated between local training and training on the RSU. This optimization aims to enhance resource utilization and task completion efficiency. Next, we will delve into the specific methods of task allocation.
@@
$N_{t}^{n}=\frac{T-t^{max}}{T_{t,L}^{n,comp}}$ (15)
where $t^{max}$ represents the maximum time used for data transmission. Thus, $T-t^{max}$ indicates the total time available for computing within a slot. In practical, the CPU typically handles multiple tasks simultaneously to improve CPU utilization. These tasks consume some CPU computing resources, resulting in the actual time available for local iterations being less than $T$. We denote the true training time as $T^{\prime}$. Therefore, the actual total local training iterations ${N_{t}^{n}}^{\ast}$ can be calculated as
${N_{t}^{n}}^{\ast}=\frac{T^{\prime}-t^{max}}{T_{t,L}^{n,comp}}$ (16)
Substitute Eq. (4) into Eq. (15) and Eq. (16), the final results can be obtained, respectively.
$N_{t}^{n}=\frac{(T-t^{max})f_{t,L}^{n,comp}}{cZ}$ (17)
${N_{t}^{n}}^{\ast}=\frac{(T^{\prime}-t^{max})f_{t,L}^{n,comp}}{cZ}$ (18)
In this paper, we study a binary partial offloading problem, where a binary variable $g_{t}^{n}$ is used to indicate whether vehicle $n$ needs to offload partial task to the RSU. $g_{t}^{n}$ is related to the assignment ratio $q_{t}^{n}$ and the assignment ratio threshold $q_{0}$. When $q_{t}^{n}$ is less than the assignment ratio threshold $q_{0}$, it indicates that vehicle $n$ will not offload partial task to RSU and reset $q_{t}^{n}=0$; otherwise, it will offload it. Therefore, $g_{t}^{n}$ can be expressed as
$g_{t}^{n}=\begin{cases}1,&\text{if }q_{t}^{n}\geq q_{0}\\ 0,&\text{if }q_{t}^{n}<q_{0}\end{cases}$ (19)
@@
We denote the total task that vehicle $n$ needs to handle in slot $t$ as $N_{t}^{n,total}$, which includes the sum of the total training iterations $N_{t}^{n}$ in slot $t$ and the remaining unprocessed training iterations $B_{t-1}^{n}$ from the previous slot. When vehicle $n$ offloads partial task to RSU, the total task $N_{t}^{n,total}$ will be divided into local training iterations $N_{t,L}^{n}$ and offloaded training iterations $N_{t,R}^{n,off}$ according to the allocated assignment ratio $q_{t}^{n}$. Specifically, RSU first determines the total training iterations $N_{R}$ in RSU side as follows
$N_{R}=\frac{T-t^{max}-T_{t,R}^{n,trans}}{T_{t,R}^{comp}}$ (20)
Thus, the expected offloaded training iterations $N_{t,R}^{n,off}$ for RSU can be calculated as
$N_{t,R}^{n,off}=g_{t}^{n}q_{t}^{n}N_{R}$ (21)
@@
$N_{t,R}^{n,off}\leftarrow\min\left(N_{t,R}^{n,off},N_{t}^{n,total}\right)$ (22)
@@
$B_{t}^{n}=\begin{cases}0,&\text{if }N_{t}^{n,re}\leq{N_{t}^{n}}^{\ast}\\ N_{t}^{n,re}-{N_{t}^{n}}^{\ast},&\text{if }N_{t}^{n,re}>{N_{t}^{n}}^{\ast}\end{cases}$ (25)
@@
Increasing the number of local iterations typically enhances classification accuracy, but it also leads to increased energy consumption per round of FL. To find a balance between classification accuracy and energy consumption, we shall formulate an optimization model to find its solution.
@@
To achieve efficient task offloading and resource allocation, it is essential for BS to consider the long-term impact of their actions during the decision-making process. Therefore, we aim to minimize the sum of the energy consumption of all vehicles in each slot. Let =$\mathbb{P}=\left\{p_{t,R}^{1},p_{t,R}^{2},\cdots p_{t,R}^{n},\cdots p_{t,R}^{N}\right\}$, $\mathbb{F}=\left\{f_{t,L}^{1,comp},f_{t,L}^{2,comp},\cdots f_{t,L}^{n,comp},\cdots f_{t,L}^{N,comp}\right\}$ and =$\mathbb{Q}=\left\{q_{t}^{1},q_{t}^{2},\cdots q_{t}^{n},\cdots q_{t}^{N}\right\}$. Then, we can formulate the optimization problem in slot $t$ as follows
@@
$g_{t}^{n}\in\left\{0,1\right\},\forall n\in[1,N]$ (29a)
$p^{min}\leq p_{t,R}^{n}\leq p^{max},\forall n\in[1,N]$ (29b)
$f^{min}\leq f_{t,L}^{n,comp}\leq f^{max},\forall n\in[1,N]$ (29c)
$0\leq q_{t}^{n}\leq 1,\forall n\in[1,N]$ (29d)
@@
After the BS allocates the assignment ratio $q_{t}^{n}$ for vehicle $n$ in slot $t$, the iterations number of offloading task $N_{t}^{n,off}$ assigned to vehicle $n$ might exceed the total number of tasks $N_{t}^{n,total}$ that the vehicle $n$ needs to process in slot $t$ [28]. We denote the excess part as the overload $O_{t}^{n}$, which can be calculated as
$O_{t}^{n}=\max\left(0,N_{t,R}^{n,off}-N_{t}^{n,total}\right)$ (31)
@@
$\mathcal{R}_{0}=\frac{1}{NS^{max}}\left(\sum_{n=1}^{N}\sum_{t=1}^{S^{max}}{\frac{O_{t}^{n}}{N_{t}^{n,total}}}\right)$ (32)